# Unbounded entropy production and violent fragmentation in long-range interacting super-Tonks-Girardeau systems

### v1

@@ Abstract
We study the non-equilibrium dynamics of a one-dimensional Bose gas with long-range interactions that decay as $(\frac{1}{r^{\alpha}})$ $(0.5<\alpha<4.0$). We investigate the system when the interactions are suddenly switched from the Tonks-Girardeau (TG) to the super-Tonks-Girardeau (sTG) limit. We find that relaxation is achieved through a complex intermediate dynamics demonstrated by violent fragmentation and chaotic delocalization. We establish that the sTG gas exhibits classical gaseous characteristics and an asymptotic state associated with unbounded entropy production. The phase diagram shows an exponential boundary between the nonthermal (quantum) gas and the thermal (classical) gas. We show the universality of the dynamics by also presenting results for spinless fermions. Weaker quench protocols give a certain degree of control over the relaxation process and induce some intermediate quasi-prethermal states. Our study showcases the complex relaxation behavior of tunable long-range interacting systems that could be engineered in state-of-the-art experiments, e.g. in trapped ions or Rydberg atoms.
entropy, long-range interactions, crystal states, super-Tonks-Girardeau quenches
I Introduction
@@
The dynamical properties of long-range interacting systems significantly differ from those of short-range interacting ones in many aspects [12, 13, 14, 15]. Depending on the exponent $\alpha$, long-range interactions can exhibit novel features. This is especially true in the very long-ranged regime when $\alpha<d$ ($d$ being the dimension of the system). These include long-lived metastable states [16, 17], supersolids and dipolar supersolids [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], dynamical phenomena such as time crystals [29, 30, 31, 32] and various Floquet phases [33, 34, 35, 36, 37].
The nonequilibrium dynamics and (pre)thermalization of isolated quantum systems with short-range interactions is rather well understood and has been extensively studied [38, 39, 40, 41, 42, 43, 40]. On the other hand, long-range interactions introduce more intriguing complexities which are poorly understood [44, 45, 46, 1]. For example, short-range interacting systems typically give rise to a single time scale of thermalization, while for strong long-range interactions thermalization can occur in multiple time steps [46].
@@
The simplest protocol for probing such nonequilibrium dynamics is a quantum quench, which consists of preparing the initial setup in the ground state of a given Hamiltonian and then suddenly quench one of its parameters to a different value at time zero. Thermalization [47] happens when a local observable approaches a statistical value or equilibrium [48, 49]. However, depending on the quench parameter, an initial quantum system can exhibit a two step relaxation process instead. The local observable first settles to a quasi-stationary value, and only at a later stage reaches thermalization. This phenomenon is termed prethermalization [50]. Prethermalization happens when there is a clear separation in relevant time scales with different physical origins, which typically depend on individual systems [51].
Prethermalization was directly observed in a one-dimensional (1D) Bose gas of 87Rb atoms [52]. In this experiment, the initial-phase fluctuating 1D Bose gas is split into two uncoupled gases in a double well potential with almost identical phase difference. The phase coherence between the two gases is further studied as a function of the evolution time. After some initial rapid evolution, a thermal-like steady state is approached – which is associated with prethermalization – before the system eventually equilibrates.
@@
Exact theoretical studies of (pre)thermalization in long-range interacting systems for both bosons and fermions are very limited [53, 13]. Study on the quench dynamics with long-range interacting spin chains so far concluded that prehermalization occurs for $10<\alpha<$ 1 [46]. In another theoretical study, utilizing spatial inhomogeneity, three different relaxation processes were observed: prethermalization only, prethermalization followed by thermalization, and thermalization only [54]. Prethermalization has also captured the attention in spinless fermions and in the context of fermionic Hubbard models [55, 56, 57]. Although significant amount of work both theoretical and experimental has been devoted to understanding the prethermalization process, its physical origin is still elusive. It is also not clear whether it can be related to the quasi-integrability of the system. For bosons, theoretical studies so far addressed the cases of quasi-BEC regimes [51] and the case of strongly interacting dipolar bosons remains rather unexplored.
@@
Our main motivation is to establish an all-in-one out-of-equilibrium dynamics for the generalized long-range interactions that can be observed in AMO experiments. To that end, the present work focuses on generalized long-range interacting bosons and fermions in the strongly interacting regime given by the Tonks-Girardeau (TG) limit. The TG gas is one of the most remarkable examples of quantum correlation amplification arising from a strong interplay between interatomic interactions and one-dimensional (1D) confinement. In the TG limit, bosons with strong repulsive interactions escape their spatial overlap to minimize their interaction energy and thereby acquire fermionic properties [58], which have been experimentally observed [59]. From a TG gas, it is possible to stabilize another highly excited and correlated state by quenching the interactions to the strongly attractive regime instead. This state is called a super-Tonks-Girardeau gas (sTG) [60, 61]. Intuitively, strong attractions should lead the system to quickly cluster. Remarkably, instead, the TG arrangement is preserved and the many-body phase relaxes to a metastable state, as verified experimentally in the work of Haller et al. [62]. In the experiment, a magnetically induced Feshbach resonance [63] is employed to control the 1D scattering properties by tuning the coupling constant of the 1D $\delta$-function contact potential $(g_{1D})$ [64]. The TG to sTG quench is achieved when $g_{1D}$ is switched from +$\infty$+ ∞ to $\infty$- ∞, which leads to the excited gas phase.
Theoretically, the sTG quench problem has been addressed by solving the dynamics of a Lieb-Liniger Bose gas [65], but paradigmatic questions regarding relaxation dynamics are still not clarified: i) What is the underlying microscopic dynamics that transforms the TG phase into the sTG phase? ii) How does the non equilibrium dynamics change when longer-ranged interactions are employed? iii) How can the excited gas phase be characterized from a many-body perspective? iv) Does equilibration happen in a single or dual time scale? v) Is it possible to control the relaxation process in the quench dynamics? Motivated by the recent experiments with magnetic atoms [66, 67], polar molecules [68, 69] and the flexibility to tune the long-range interactions [70], we address all these questions by studying sTG quenches for long-ranged systems with varying exponent $\alpha$.
We solve the time-dependent Schrödinger equation (TDSE) for the sTG quench process numerically by employing the MultiConfiguration Time-Dependent Hartree method for bosons (MCTDHB) [71, 72, 73, 74] implemented in the software MCTDH-X [74, 75, 76, 77, 78, 79, 80]. MCTDH-X has been a very prolific tool for studying the dynamics of long-range interacting quantum gases in the past [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98], as it enables us to evaluate the full time evolution of time-dependent processes in a very efficient way.
We follow the experimental protocol of Haller [62] and consider generic power-law interactions with $40.5\leq\alpha<$ 4.
@@
For each value of $\alpha$, we first prepare $N=5$ strongly repulsive bosons in the ground state of a harmonic trap, obtaining a crystal state. Then, at time $t$=0, we suddenly quench their interactions to the strongly attractive regime and monitor the particle dynamics. We find that the time evolution exhibits violent dynamical fragmentation, unbounded entropy production, and complete delocalization in configuration space. These three features suggest that the highly excited sTG phase behaves as a classical gas when starting from a crystal state exhibiting long-range interactions. We also perform analogous simulations with fermions, observing a very similar behavior which indicates the universality of our results in the dynamics of non-local interacting systems. For both bosons and fermions, the system eventually relaxes to a fully occupied configuration space for the entire range of $\alpha$. We do not find any signatures of any intermediate time scale in the relaxation process. However, we observe an exponential boundary in the entropy evolution between relaxed and non-relaxed states, indicating a progressive slowdown in the relaxation process at increasing longer-ranged interactions (lower value of $\alpha$). To understand the process of relaxation, the system is also studied for quenches to other weakly attractive regimes. We observe that the relaxation process is less violent, slows down and begins to exhibit very short lived quasi-equilibrium states when $\alpha<1$.
@@
The paper is structured as follows. In Sec. II, we discuss the Hamiltonian and theoretical approach. In Sec. III, we introduce quantities of interest. Sec. IV discusses the quantum to classical behavior through the analysis of three key quantities: orbital occupation, Fock space occupation, and entropy measures. Sec. V illustrates the underlying relaxation process both for bosons and fermions. Sec. VI presents how to control the quench mechanism to achieve almost prethermal states. We conclude our paper in Sec. VII. Appendix A presents an overview of the initial states and Appendix B considers a stronger quench dynamics.
II Model and theoretical approach
@@
The Hamiltonian $\hat{H}$ is expressed in dimensionless units  [^1] The Hamiltonian $\hat{H}$ is obtained by dividing the dimensionful Hamiltonian by $\frac{\hbar^{2}}{mL^{2}}$, with $m$ the mass of the particles and $L$ an arbitrary length scale, which we choose to be the point at which the harmonic trap equals $\frac{1}{2}$.. The operator $\hat{h}(x)=\hat{T}(x)+\hat{V}(x)$ is the one-body Hamiltonian containing the kinetic energy $\hat{T}(x)=-\frac{1}{2}\hat{\partial}_{x}^{2}$ and the external potential $\hat{V}(x)=\frac{1}{2}x^{2}$. The operator $\hat{W}(x_{i}-x_{j})$ describes the two-body interaction between particles at positions $x_{i}$ and $x_{j}$ (which can be either attractive or repulsive).
The MCTDH ansatz for the bosonic many-body wave function is a linear combination of time dependent permanents constructed over $M$ single-particle wave functions called orbitals  [^2] The ansatz for fermionic many-body wave functions is equivalent, but Slater determinants replace the permanents. We redirect to Ref. [75] for further details.:
@@
It is important to emphasize that in the ansatz (3), both the expansion coefficients $C_{\vec{n}}(t)|\vec{n};t\rangle$ and the orbitals that build up the permanents $|\vec{n};t\rangle$ are time-dependent, fully variationally optimized quantities. This time-adaptive basis allows the sampled Hilbert space to dynamically follow the motion under any quench dynamics.
In the limit $M\rightarrow\infty$, the representation of Eq. 3 is exact. However, in practice, the size of the Hilbert space has to be truncated by using a finite value of $M$ during computation. The time-dependent permanents assure that a given degree of accuracy is reached with a much shorter expansion compared to the time independent basis used in exact diagonalization.
@@
as a spectral decomposition. In the expression for the reduced one-body density matrix, $\hat{\psi}^{(\dagger)}(x)$ are field operators for the creation and annihilation of one particle at position $x$. For a non-fragmented or condensed state, the occupation of the lowest natural orbital is close to unity. However, for a fragmented state, several natural orbitals may contribute significantly. In particular, for bosons, (lack of) fragmentation illustrates how close the many-body state is to a global superfluid that can be described by a single wave function.
III Quantities of interest
To study statistical relaxation, the Shannon information entropy (SIE) $S^{\mathrm{info}}(t)$ is the ideal quantity [101]. The SIE of the one-body density in coordinate space is defined as $S^{\mathrm{info}}(t)=-\int{\mathrm{d}x\rho^{(1)}(x,t)\ln\left[{\rho^{(1)}(x,t)}\right]}$. The SIE is a measure of localization/delocalization of the corresponding density. However, it is insensitive to the correlation present in $|\psi(t)\rangle$. Considering the MCTDHB ansatz (3), then, we define a few alternative entropy measures [83]. The coefficient Shannon information entropy is defined as
@@ III Quantities of interest
It characterizes the distribution of the state $|\psi(t)\rangle$ in the Fock space. $S_{C}=0$ indicates a single-configuration state or a mean-field state. When $|\psi(t)\rangle$ spreads over the available space determined by the configurations, several expansion coefficients become nonzero and $S_{C}$ gradually increases. When the $N$-body Fock space is completely populated, $S_{C}$ saturates to the maximal value. Thus, in a quench procedure, $S_{C}(t)$ measures the degree of Fock space delocalization with time.
The occupation Shannon information entropy is defined as
@@
where $\bar{n}_{i}(t)=\frac{n_{i}(t)}{N}$ are natural occupations normalized with the particle number. In the single-orbital mean-field theory, as the reduced density matrix has only a single eigenvalue, $S_{n}=0$ and the corresponding many-body state is non-fragmented. When multiple significant orbitals contribute, the corresponding many-body state becomes fragmented. Thus, in a relaxation process, $S_{n}(t)$ initially starts with a minimum value (which may not be zero), increases and finally saturates to a maximal value. Therefore, $S_{n}(t)$ conveys the emergence and degree of fragmentation.
Finally, we can also study entropy measures obtained from the coefficients of the many-body decomposition. The $N$-body coefficient entropy $S_{C}^{N}(t)$ is defined as a hybrid version of the coefficient Shannon information entropy,
This quantity conveys information about delocalization in Fock space, but considering all possible pairings of configurations. A related quantity is the coefficient inverse participation ratio. This is another measure to detect the effective number of basis states participating in the time evolution of the many-body state and is generally utilized to understand irregular dynamics. It is defined as
$I_{C}(t)=\frac{1}{\sum_{\vec{n}}|C_{\vec{n}}(t)|^{4}}.$ (9)
When $I_{C}(t)$ is small, the corresponding many-body state is close to the mean-field state and the corresponding dynamics is regular. Whereas large $I_{C}(t)$ indicates that the corresponding dynamics is irregular or chaotic.
IV Quantum to classical behavior

### v2

# Unbounded entropy production and violent fragmentation for repulsive-to-attractive interaction quench in long-range interacting systems

@@ Abstract
We study the non-equilibrium dynamics of a one-dimensional Bose gas with long-range interactions that decay as $(\frac{1}{r^{\alpha}})$ $(0.5<\alpha<4.0$). We investigate exotic dynamics when the interactions are suddenly switched from strongly repulsive to strongly attractive, a procedure known to generate super-Tonks-Girardeau gases in systems with contact interactions. We find that relaxation is achieved through a complex intermediate dynamics demonstrated by violent fragmentation and chaotic delocalization. We establish that the relaxed state exhibits classical gaseous characteristics and an asymptotic state associated with unbounded entropy production. The phase diagram shows an exponential boundary between the coherent (quantum) gas and the chaotic (classical) gas. We show the universality of the dynamics by also presenting analogous results for spinless fermions. Weaker quench protocols give a certain degree of control over the relaxation process and induce a slower initial entropy growth. Our study showcases the complex relaxation behavior of tunable long-range interacting systems that could be engineered in state-of-the-art experiments, e.g. in trapped ions or Rydberg atoms.
1 Introduction
@@
The dynamical properties of long-range interacting systems significantly differ from those of short-range interacting ones in many aspects [12, 13, 14, 15]. Depending on the exponent $\alpha$, long-range interactions can exhibit novel features. This is especially true in the very long-ranged regime when $\alpha<d$ ($d$ being the dimension of the system). These include long-lived metastable states [16, 17], supersolids and dipolar supersolids [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], dynamical phenomena such as time crystals [29, 30, 31, 32] and various Floquet phases [33, 34, 35, 36, 37]. The nonequilibrium dynamics and relaxation process of isolated quantum systems with short-range interactions is rather well understood and has been extensively studied [38, 39, 40, 41, 42, 43, 40]. On the other hand, long-range interactions introduce more intriguing complexities which are poorly understood [44, 45, 46, 1]. For example, short-range interacting systems typically give rise to a single time scale of relaxation, while for strong long-range interactions relaxation can occur in multiple time steps [46].
@@
The simplest protocol for probing exotic nonequilibrium dynamics is a quantum quench, which consists of preparing the initial setup in the ground state of a given Hamiltonian and then suddenly switch one of its parameters to a different value at time zero. A quantum quench allows to probe the fate of a quantum system when perturbed away from equilibrium. Thermalization [47] happens when a local observable approaches a statistical value or equilibrium [48, 49]. However, depending on the quench parameter, an initial quantum system can exhibit a different relaxation process instead. One example is a phenomenon termed prethermalization [50], where the local observable first settles to a quasi-stationary value, and only at a later stage reaches thermalization. Prethermalization happens when there is a clear separation in relevant time scales with different physical origins, which typically depend on individual systems [51, 52].
@@
Exact theoretical studies on the relaxation pathway in long-range interacting systems for both bosons and fermions are very limited [53, 13]. Studies on the quench dynamics with long-range interacting spin chains so far concluded that prethermalization occurs for $10<\alpha<$ 1 [46]. In another theoretical study, utilizing spatial inhomogeneity, three different relaxation processes were observed: prethermalization only, prethermalization followed by thermalization, and thermalization only [54]. Prethermalization has also captured the attention in spinless fermions and in the context of fermionic Hubbard models [55, 56, 57]. Although significant amount of work both theoretical and experimental has been devoted to understanding the prethermalization process, its physical origin is still elusive. It is also not clear whether it can be related to the quasi-integrability of the system. For bosons, theoretical studies so far addressed the cases of quasi-BEC regimes [51] and the case of strongly interacting dipolar bosons remains rather unexplored.
@@
Our main motivation is to establish an all-in-one out-of-equilibrium dynamics for the generalized long-range interactions that can be observed in AMO experiments with magnetic atoms [58, 59] and polar molecules [60, 61]. To that end, the present work focuses on generalized long-range interacting bosons and fermions where the interactions are quenched from the strongly repulsive to the strongly attractive regime. This procedure is akin to the approaches used to study the stability of strongly interacting gases with contact interactions starting from the Tonks-Girardeau (TG) regime. The strongly interacting bosons in TG limit exhibits remarkable effect of quantum correlation [62]. Due to strong repulsive interactions, atoms escape their spatial overlap to minimize their interaction energy and thereby acquire fermionic properties [62], which have been experimentally observed [63]. By quenching the interactions to the strongly attractive regime instead makes a transition to another highly excited and correlated state, called a super-Tonks-Girardeau gas (sTG) [64, 65] which has been verified experimentally in the work of Haller et al. [66]. The nature of this highly correlated gas can be characterized by either breathing dynamics or confirmation that it neither clusters, nor remains bound, but it is metastable [67, 68]. Theoretically, quench problems like the ones described above have been addressed by solving the dynamics of a Lieb-Liniger Bose gas [69], but several paradigmatic questions regarding relaxation dynamics are still not clarified.
In this work, we follow the experimental protocol of Haller [66] and consider generic power-law interactions with $40.5\leq\alpha<$ 4. For each value of $\alpha$, we first prepare $N=5$ strongly repulsive bosons in the ground state of a harmonic trap, obtaining a crystal state. Then, at time $t$=0, we suddenly quench their interactions to the strongly attractive regime and monitor the out-of-equlibrium dynamics. We solve the time-dependent Schrödinger equation (TDSE) for the quench process numerically by employing the MultiConfiguration Time-Dependent Hartree method for bosons (MCTDHB) [70, 71, 72, 73]. The many-body dynamics and the relaxation process are further studied by the evolution of several measures of many-body information entropy which answer the following questions: i) What is the underlying microscopic dynamics that transforms the strongly repulsive phase into the strongly attractive one? ii) How does the non equilibrium dynamics change when longer-ranged interactions are employed? iii) How can the excited gas phase be characterized from a many-body perspective? iv) Does equilibration happen in a single or dual time scale? v) Is it possible to control the relaxation process in the quench dynamics? Tuning the long-range interactions [74], we address all these questions by studying quenches for long-ranged systems with varying exponent $\alpha$. To demonstrate the universality of the features we observe, we also perform analogous quench procedures on fermions, by employing the fermionic variant of the MCTDH method (MCTDHF). Both methods are implemented in the software MCTDH-X [73, 75, 76, 77, 78, 79, 80]. MCTDH-X has been a very prolific tool for studying the dynamics of long-range interacting quantum gases in the past [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98], as it enables us to evaluate the full time evolution of time-dependent processes in a very efficient way.
@@
In our simulations, we find that the time evolution exhibits violent dynamical fragmentation, unbounded entropy production, and complete delocalization in configuration space. These three features suggest that the highly excited post-quench phase behaves as a quasiclassical gas when starting from a crystal state exhibiting long-range interactions. From the analogous simulations performed on fermions, we observe a very similar behavior which indicates the universality of our results in the dynamics of non-local interacting systems. For both bosons and fermions we observe very exotic dynamics although the system eventually relaxes to a fully occupied configuration space for the entire range of $\alpha$. Somewhat surprisingly, we do not find any signatures of any intermediate time scale in the relaxation process. However, we observe an exponential boundary in the entropy evolution between relaxed and non-relaxed states, indicating a progressive slowdown in the relaxation process at increasing longer-ranged interactions (lower value of $\alpha$). To understand the process of relaxation, the system is also studied for quenches to other weakly attractive regimes. We observe that the relaxation process is less violent, slows down and begins to exhibit a quadratic entropy scaling with a very slow growth at short times when $\alpha<1$.
@@
The paper is structured as follows. In Sec. II, we discuss the Hamiltonian and theoretical approach. In Sec. III, we introduce quantities of interest. Sec. IV discusses the coherent to chaotic behavior through the analysis of three key quantities: orbital occupation, Fock space occupation, and entropy measures. Sec. V illustrates the underlying relaxation process both for bosons and fermions. Sec. VI presents how to control the quench mechanism to achieve a drastic dynamics slowdown. We conclude our paper in Sec VII. Appendix A presents an overview of the initial states, Appendix B considers a stronger quench dynamics, and Appendix C summarizes the units used in the numerics.
2 Model and theoretical approach
@@
The Hamiltonian $\hat{H}$ is expressed in dimensionless units Please refer to appendix C for a complete discussion of how to obtain dimensionless units in MCTDH-X  [^2] The Hamiltonian $\hat{H}$ is obtained by dividing the dimensionful Hamiltonian by $\frac{\hbar^{2}}{mL^{2}}$, with $m$ the mass of the particles and $L$ an arbitrary length scale, which we choose to be the point at which the harmonic trap equals $\frac{1}{2}$.. The operator $\hat{h}(x)=\hat{T}(x)+\hat{V}(x)$ is the one-body Hamiltonian containing the kinetic energy $\hat{T}(x)=-\frac{1}{2}\hat{\partial}_{x}^{2}$ and the external potential $\hat{V}(x)=\frac{1}{2}x^{2}$. The operator $\hat{W}(x_{i}-x_{j})$ describes the two-body interaction between particles at positions $x_{i}$ and $x_{j}$ (which can be either attractive or repulsive).
The MCTDH ansatz for the bosonic many-body wave function is a linear combination of time dependent permanents constructed over $M$ single-particle wave functions called orbitals. The ansatz for the fermionic many-body wave function is equivalent, but Slater determinants replace the permanents. We redirect to Ref. [75] for further details. Both cases can be written as
@@
It is important to emphasize that in the ansatz (3), both the expansion coefficients $C_{\vec{n}}(t)|\vec{n};t\rangle$ and the orbitals that build up the configurations $|\vec{n};t\rangle$ are time-dependent, fully variationally optimized quantities. This time-adaptive basis allows the sampled Hilbert space to dynamically follow the motion under any quench dynamics.
In the limit $M\rightarrow\infty$, the representation of Eq. 3 is exact. However, in practice, the size of the Hilbert space has to be truncated by using a finite value of $M$ during computation. The time-dependent orbitals assure that a given degree of accuracy is reached with a much shorter expansion compared to the time independent basis used in exact diagonalization.
@@
as a spectral decomposition. In the expression for the reduced one-body density matrix, $\hat{\psi}^{\dagger}(x)$ and $\hat{\psi}(x)$ are field operators for the creation and annihilation of one particle at position $x$. For a non-fragmented or condensed state, the occupation of the lowest natural orbital is close to unity. However, for a fragmented state, several natural orbitals may contribute significantly. In particular, for bosons, (lack of) fragmentation illustrates how close the many-body state is to a global superfluid that can be described by a single wave function.
3 Quantities of interest
To study statistical relaxation, the Shannon information entropy (SIE) $S^{\mathrm{info}}(t)$ is the ideal quantity [99]. The SIE of the one-body density in coordinate space is defined as $S^{\mathrm{info}}(t)=-\int{\mathrm{d}x\rho^{(1)}(x,t)\ln\left[{\rho^{(1)}(x,t)}\right]}$. The SIE is a measure of localization/delocalization of the corresponding density. However, it is insensitive to the correlation present in $|\psi(t)\rangle$. Considering the MCTDH ansatz (3), then, we define a few alternative entropy measures [83].
We begin by considering information that can be extracted from the coefficients of the MCTDH expansion. The modulus squared of each coefficient can be expressed as
$|C_{\vec{n}}(t)|^{2}=\frac{1}{\prod_{i=1}^{M}n_{i}!}\left<\Psi\right|(\hat{b}_{1}(t))^{n_{1}}\cdots(\hat{b}_{M}(t))^{n_{m}}(\hat{b}_{1}^{\dagger}(t))^{n_{1}}\cdots(\hat{b}_{M}^{\dagger}(t))^{n_{m}}\left|\Psi\right>$ (6)
indicating that – depending on the state – one, some, or all $M$ creation and annihilation operators may contribute to the value of the coefficient $|C_{\vec{n}}(t)|^{2}$. Consequently, the distribution of these coefficients provides a direct qualitative assessment of the many-body entropies within the system.
The coefficient Shannon information entropy is defined as
@@ 3 Quantities of interest
It characterizes the distribution of the state $|\psi(t)\rangle$ in the Fock space. A related quantity is the coefficient inverse participation ratio, defined as
$I_{C}(t)=\frac{1}{\sum_{\vec{n}}|C_{\vec{n}}(t)|^{4}}.$ (8)
This is another measure to detect the effective number of basis states participating in the time evolution of the many-body state and is generally utilized to understand irregular dynamics. We can also define a hybrid version of the coefficient Shannon information entropy, the $N$-body coefficient entropy $S_{C}^{N}(t)$, as
This quantity also conveys information about delocalization in Fock space, but considering all possible pairings of configurations. In terms of properties, it inherits most of its features from the coefficient entropy $S_{C}(t)$. Therefore, in the following, we will focus our discussion on $S_{C}(t)$ and $I_{C}(t)$.
We highlight a few limiting cases for these two quantities. When the system is (and remains) in a single-configuration state, $S_{C}(t)=0$ and $I_{C}(t)=1$. This type of state is equivalent to a mean-field configuration. By definition, then, mean-field approaches such as the Gross-Pitaevskii equation or multiorbital methods will always lead to a trivial coefficient entropy and inverse participation ratio. In MCTDHB, instead, these quantities can acquire much larger values.
At the other extreme, when the $N$-body Fock space is completely and uniformly populated, $S_{C}(t)$ and $I_{C}(t)$ will saturate to their maximal values. This situation indicates that the MCTDHB expansion has exhausted all the available configuration space to describe the many-body state. This is typical of systems with strong interactions or significant entanglement, where the particles do not behave independently but are instead highly correlated. Another example is chaotic systems which sample the entire configuration space. Strictly speaking, this situation can occur only in the limit $M\to\infty$. If $S_{C}(t)$ and $I_{C}(t)$ show an unbound increase when evaluated for progressively larger configuration space (larger values of $M$), it is an indication of chaotic-like or quasi-classical (incoherent) behavior.
At intermediate values of $S_{C}(t)$ and $I_{C}(t)$, $|\psi(t)\rangle$ spreads over only a part of the available configuration space and several but not all expansion coefficients are nonzero. In this case, the coefficient entropy gives a quantitative measure of the many-body character of the state at time $t$, which interpolates between the fully coherent, single-configuration case, and the fully chaotic, infinite-configuration case. Similarly, when $I_{C}(t)$ is small, the corresponding many-body state is close to the mean-field state and the corresponding dynamics is regular. Large $I_{C}(t)$ indicates instead that the corresponding dynamics is irregular or chaotic.
In a quench procedure like the one we study in this work, we typically start with a moderately correlated state exhibiting a low or intermediate value of $S_{C}(t)$ and $I_{C}(t)$, and we then observe their time evolution. If a lot of energy is injected into the system, $S_{C}(t)$ and $I_{C}(t)$ will typically increase in time, indicating spread of correlations, Fock-space delocalization, and progress towards a chaotic state.
While the entropies introduced above give a good grasp of the many-body character of the state and its potential progress towards a chaotic configuration, they are based on coefficients alone, which are not uniquely defined and instead depend on the many-body state decomposition into orbitals (i.e. the permanents). Rotating the orbitals by a unitary transformation can be compensated by transforming the coefficients. Thus, any MCTDH wavefunction has infinite representations: $\left|\Psi\right>=\sum_{\vec{n}}C_{\vec{n}}\left|n\right>=\sum_{\vec{m}}C_{\vec{m}}\left|m\right>$ etc, i.e. the same many-particle wavefunction can thus be described by different orbitals and different coefficients. While the MCTDH variational procedure converges to a state with the most optimized natural occupation, and this correlates with a restricted configuration population, there is no guarantee that the achieved configuration will be “minimal” in some sense. There is a way of selecting coefficients with the narrowest distribution based on variance calculations [70], but they are quite cumbersome and deserve a standalone treatment, which is beyond the scope of our study.
To obviate to the lack of invariance for the coefficients, we thus consider another measure of entropy obtained from explicitly invariant quantities, namely the natural occupations defined in Eq. 5. This quantity is called occupation Shannon information entropy and is defined as
@@
where $\bar{n}_{i}(t)=\frac{n_{i}(t)}{N}$ are natural occupations normalized with the particle number.
This entropy quantifies the distribution of particles among the various natural orbitals. A lower occupation entropy indicates that particles are predominantly occupying fewer orbitals (signifying higher coherence), whereas a higher entropy suggests a more uniform distribution across multiple orbitals (indicative of fragmentation and reduced coherence).
The limiting cases for this quantity are similar to those of the coefficient entropy. In a single-orbital mean-field theory, as the reduced density matrix has only a single eigenvalue, $S_{n}(t)=0$ and the corresponding many-body state is non-fragmented. At the opposite extreme, when $M\to\infty$ and all the orbitals have a finite (uniform) occupation, we are in a completely chaotic regime. In the intermediate cases, when multiple significant orbitals contribute but the occupation is not uniform at $1/M$, the corresponding many-body state is fragmented but retains a certain degree of coherence.
In a relaxation process, $S_{n}(t)$ initially starts with a minimum value (which may not be zero), increases, and finally saturates to a maximal value. Therefore, $S_{n}(t)$ conveys the emergence and degree of fragmentation over time. If $S_{n}(t)$ appears to increase unboundedly when the same simulation (same initial conditions) is performed but with an increasing number of orbitals, it is a sign that the system relaxes towards a chaotic state where all occupations are macroscopic, i.e. all the eigenvalues of the reduced one-body density matrix have the same magnitude.
We remark that even for multiorbital mean-field theories, which can exhibit a finite value of $S_{n}(t)$, this value stays constant throughout the dynamics because the populations cannot change. Thus, this quantity is a prerogative of multiconfigurational methods like MCTDHB only.
4 Quantum to classical behavior